# Formation of millisecond pulsar-helium star binaries

### v1

@@ 3 Results
Figure 2 presents the initial companion mass-orbital period distribution for NS X-ray binaries, alongside their corresponding final evolutionary outcomes. Here, we set CE efficiency $\alpha_{\rm CE}=1$, initial NS magnetic field $B_{0}=10^{12}$ G, and initial spin period $P_{\rm s,0}=1$ s as fiducial parameters. The figure shows that systems with initial orbital periods exceeding $\sim$1000 days typically do not undergo RLO MT (triangles). These wide binaries may evolve into symbiotic binaries when the companion star becomes a red giant star or asymptotic giant star (Lü et al., 2012; Yungelson et al., 2019; Deng & Li, 2024a). Only part of the regions with initial companion masses $\leq 4.5\,M_{\sun}$ undergo stable MT (stars). These systems can typically evolve into partially recycled pulsar binaries (Tauris et al., 2000; Shao & Li, 2012). Apart from this, all other regions (squares, diamonds, and circles) undergo unstable MT and enter a CE evolution phase. Squares and diamonds represent systems undergoing unstable Case B and Case C mass transfer, respectively, both leading to the formation of NS + He star binaries after successful CE ejection. They can eventually evolve into NS-WD binaries or DNSs (provided the second supernova does not disrupt the binary; Shao & Li, 2018; Vigna-Gómez et al., 2018; Deng et al., 2024). On the other hand, if the binary systems fail to eject the envelope during the CE phase, the binaries merge (circles), potentially giving rise to Thorne-Źytkow objects (Thorne, 1974; Podsiadlowski et al., 1995).
@@
Focusing on NS + He star progenitor systems, depicted by the squares and diamonds in Figure 2, we find that the parameter space for successful CE ejection diminishes with increasing initial companion star mass (see also, Gallegos-Garcia et al., 2023; Nie et al., 2025). Only two regions (red squares) yield NS + He star systems with the NS spin periods below 10 ms. The parameter space for the formation of NSs with spin periods between 103010-30 ms, 3010030-100 ms, and 1001000100-1000 ms progressively expands. This implies that, in case of super-Eddington accretion, pre-CE RLO MT seems able to accelerate the NS spins to the millisecond regimes (see also Pan et al., 2022).
@@ 4 Discussion
The formation of NS + He star binaries encompasses several distinct accretion phases: wind accretion, RLO accretion, accretion during the CE phase, and post-CE He star wind accretion. Owing to the turbulent nature of material flow in wind accretion, the angular momentum of the captured matter may fluctuate between positive and negative values. Over extended timescales, these components largely cancel each other out. Consequently, the net torque exerted on the NS during this phase may be minimal or potentially negligible (Mao & Li, 2024). This phenomenon manifests as spin-up and spin-down reversals in wind-fed X-ray pulsars (Bildsten et al., 1997; Malacaria et al., 2020) and has been corroborated by hydrodynamical simulations (Blondin & Raymer, 2012; Xu & Stone, 2019). Based on these considerations, we disregard the influence of the wind accretion phase in this study.
@@
Similarly, accretion onto compact objects during a CE is commonly assumed to adhere to the spherically symmetric Hoyle-Lyttleton model (MacLeod & Ramirez-Ruiz, 2015; MacLeod et al., 2017; Hutchinson-Smith et al., 2024; Rosselli-Calderon et al., 2024). Under this model, a NS in the CE phase would also fail to efficiently spin up, which explains our omission of spin acceleration effects during this stage. Furthermore, if the NS had undergone a stable accretion-driven spin-up process during the CE phase, its spin axis would align with the post-CE binary orbital axis (Bhattacharya & van den Heuvel, 1991). However, in the case of PSR J1928+1815, there is a significant misalignment between the spin axis and the orbital axis (Yang et al., 2025). This feature suggests that the accretion process during the CE phase was typically unstable or directionally random (Biryukov & Abolmasov, 2021).
@@
Subsequent to the CE phase, the binary will enter a CE decoupling phase (CEDP), during which the He star may retain a residual hydrogen envelope (Nie et al., 2025). MT from this envelope onto the NS could also induce spin-up. Moreover, NS + He star systems can evolve into DNS binaries, potentially involving Case BB mass transfer (Tauris et al., 2017; Jiang et al., 2021; Deng et al., 2024; Guo et al., 2025). In summary, for NS + He or DNS systems, the initially formed NS will have undergone multiple accretion episodes, rendering the study of its spin evolution particularly intricate and challenging.
@@
Two additional potential formation channels for NS + He star binaries warrant consideration. One channel involves the double-core CE phase (Brown, 1995; Dewi et al., 2006; Andrews et al., 2015; Hwang et al., 2015). In this scenario, the initial binary mass ratio is typically near unity, potentially leading to the formation of DNS systems after the second SN, provided the system remains bound. However, the contribution of this channel to DNS formation varies significantly across different models and assumptions (Vigna-Gómez et al., 2018; Shao & Li, 2018; Deng et al., 2024). A distinctive feature of this channel is that the resulting NS + He star systems have not experienced pre-CE wind or RLO MT, nor hypercritical accretion during the CE phase. Consequently, the binary would host a non-recycled NS.
@@
Another formation channel is accretion-induced collapse (AIC) of a massive WD (see, e.g., Miyaji et al., 1980; Canal et al., 1990; Nomoto & Kondo, 1991; Freire & Tauris, 2014; Wang, 2018). In this scenario, if the compact object in an IMXB is a massive oxygen-neon-magnesium (ONeMg) WD (close to the Chandrasekhar mass, $M_{\rm Ch}$), accretion onto the WD may supply sufficient material for its mass to exceed $M_{\rm Ch}$. However, rapid rotation might prevent a super-Chandrasekhar WD from collapsing into a NS (Tauris et al., 2013b; Freire & Tauris, 2014). After the subsequent CE phase, the WD may lose angular momentum due to r-mode instability and/or magnetic dipole radiation, potentially leading to collapse into a MSP (Freire & Tauris, 2014). It should be noted that the formation of NS+He star systems through this channel requires rather stringent conditions (a narrow initial white dwarf mass range, specific accretion rate ranges, etc.), thus also introducing significant uncertainty. Further investigation is needed in the future work.
@@
The biggest uncertainty in this work pertains to CE evolution (see Ivanova et al., 2013, for a review), which determines the properties of the resulting NS + He binaries. The criteria for CE evolution, which are closely intertwined with MT stability criteria, remain a subject of ongoing debate (Soberman et al., 1997; Pavlovskii & Ivanova, 2015; Ge et al., 2020; Shao & Li, 2021; Deng et al., 2024). The post-CE evolutionary fate of a binary system depends on the relative magnitudes of the envelope binding energy and the orbital energy (Webbink, 1984; Nelemans & Tout, 2005; Klencki et al., 2021). However, significant ambiguity surrounds the choice of the core-envelope boundary for calculating the CE binding energy, and different boundary definitions can significantly impact the calculated value (Dewi & Tauris, 2000; Dominik et al., 2012; Marchant et al., 2021; Klencki et al., 2021; Gallegos-Garcia et al., 2023). The CE ejection efficiency, $\alpha_{\rm CE}$, also carries substantial uncertainty. Theoretically, $\alpha_{\rm CE}$ should be less than or equal to unity. Studies of WD binaries suggest the $\alpha_{\rm CE}$ values of 0.20.30.2-0.3 (Zorotovic et al., 2010; Davis et al., 2012; Nandez et al., 2015; Ge et al., 2022; Scherbak & Fuller, 2022; Zorotovic & Schreiber, 2022). However, research on the formation of black hole LMXBs suggests that $\alpha_{\rm CE}$ is expected to be significantly greater than unity for successful envelope ejection and binary survival  [^3] Some studies suggest that if BHs form via failed supernovae (O’Connor & Ott, 2011), the formation of BH LMXBs can be explained without invoking unrealistically high $\alpha_{\rm CE}$ values (Wang et al., 2016a; Shao & Li, 2020; Deng & Li, 2024b). (Kalogera, 1999; Podsiadlowski et al., 2003; Kiel et al., 2008; Yungelson & Lasota, 2008). Studies of Galactic DNSs also find that $\alpha_{\rm CE}$ needs to exceed unity to match observations (Chu et al., 2022; Sgalletta et al., 2023). Some researchers propose that additional, as-yet-unknown energy sources, supplementing the orbital energy, may contribute to envelope removal (Ivanova, 2002; Podsiadlowski et al., 2010). Currently, fully three-dimensional hydrodynamic simulations of the CE phase in NS HMXBs remain computationally prohibitive (e.g., Moreno et al., 2022; Vetter et al., 2024).
@@ 5 Summary
In this work, we investigate NS accretion in intermediate-mass or high-mass X-ray binaries during the RLO phase preceding CE evolution. Motivated by the discovery of nearly ten NS ULXs (Kuranov et al., 2021), we consider super-Eddington accretion-induced field decay and spin-up in the NSs. By combining MESA with the binary population synthesis code BSE, we explore the formation of NS + He star systems and determine the distributions of the orbital period, He star mass, NS spin period, magnetic field strength, and accreted mass. Our main findings are as follows:

### v2

@@ 3 Results
Figure 2 presents the initial companion mass-orbital period distribution for NS X-ray binaries, alongside their corresponding final evolutionary outcomes. Here, we set CE efficiency $\alpha_{\rm CE}=1$, initial NS magnetic field $B_{0}=10^{12}$ G, and initial spin period $P_{\rm s,0}=1$ s as fiducial parameters. The figure shows that systems with initial orbital periods exceeding $\sim$1000 days typically do not undergo RLO MT (triangles). These wide binaries may evolve into symbiotic binaries when the companion star becomes a red giant star or asymptotic giant star (Lv2012; Yungelson2019; Deng2024c). Only part of the regions with initial companion masses $\leq 4.5\,M_{\sun}$ undergo stable MT (stars). These systems can typically evolve into partially recycled pulsar binaries (Tauris2000; Shao2012). Apart from this, all other regions (squares, diamonds, and circles) undergo unstable MT and enter a CE evolution phase. Squares and diamonds represent systems undergoing unstable Case B and Case C mass transfer, respectively, both leading to the formation of NS + He star binaries after successful CE ejection. They can eventually evolve into NS-WD binaries or DNSs (provided the second supernova does not disrupt the binary; Shao2018a; Vigna2018; Deng2024a). On the other hand, if the binary systems fail to eject the envelope during the CE phase, the binaries merge (circles), potentially giving rise to Thorne-Źytkow objects (Thorne1974; Podsiadlowski1995).
@@
Focusing on NS + He star progenitor systems, depicted by the squares and diamonds in Figure 2, we find that the parameter space for successful CE ejection diminishes with increasing initial companion star mass (see also, Gallegos2023; Nie2025). Only two regions (red squares) yield NS + He star systems with the NS spin periods below 10 ms. The parameter space for the formation of NSs with spin periods between 103010-30 ms, 3010030-100 ms, and 1001000100-1000 ms progressively expands. This implies that, in case of super-Eddington accretion, pre-CE RLO MT seems able to accelerate the NS spins to the millisecond regimes (see also Pan2022).
@@ 4 Discussion
The formation of NS + He star binaries encompasses several distinct accretion phases: wind accretion, RLO accretion, accretion during the CE phase, and post-CE He star wind accretion. Owing to the turbulent nature of material flow in wind accretion, the angular momentum of the captured matter may fluctuate between positive and negative values. Over extended timescales, these components largely cancel each other out. Consequently, the net torque exerted on the NS during this phase may be minimal or potentially negligible (Mao2024). This phenomenon manifests as spin-up and spin-down reversals in wind-fed X-ray pulsars (Bildsten1997; Malacaria2020) and has been corroborated by hydrodynamical simulations (Blondin2012; Xu2019). Based on these considerations, we disregard the influence of the wind accretion phase in this study.
@@
Similarly, accretion onto compact objects during a CE is commonly assumed to adhere to the spherically symmetric Hoyle-Lyttleton model (MacLeod2015a; MacLeod2017; Hutchinson2024; Rosselli2024). Under this model, a NS in the CE phase would also fail to efficiently spin up, which explains our omission of spin acceleration effects during this stage. Furthermore, if the NS had undergone a stable accretion-driven spin-up process during the CE phase, its spin axis would align with the post-CE binary orbital axis (Bhattacharya1991). However, in the case of PSR J1928+1815, there is a significant misalignment between the spin axis and the orbital axis (Yang2025). This feature suggests that the accretion process during the CE phase was typically unstable or directionally random (Biryukov2021).
@@
Subsequent to the CE phase, the binary will enter a CE decoupling phase (CEDP), during which the He star may retain a residual hydrogen envelope (Nie2025). MT from this envelope onto the NS could also induce spin-up. Moreover, NS + He star systems can evolve into DNS binaries, potentially involving Case BB mass transfer (Tauris2017; Jiang2021; Deng2024a; Guo2025). In summary, for NS + He or DNS systems, the initially formed NS will have undergone multiple accretion episodes, rendering the study of its spin evolution particularly intricate and challenging.
@@
Two additional potential formation channels for NS + He star binaries warrant consideration. One channel involves the double-core CE phase (Brown1995; Dewi2006; Andrews2015; Hwang2015). In this scenario, the initial binary mass ratio is typically near unity, potentially leading to the formation of DNS systems after the second SN, provided the system remains bound. However, the contribution of this channel to DNS formation varies significantly across different models and assumptions (Vigna2018; Shao2018a; Deng2024a). A distinctive feature of this channel is that the resulting NS + He star systems have not experienced pre-CE wind or RLO MT, nor hypercritical accretion during the CE phase. Consequently, the binary would host a non-recycled NS.
@@
Another formation channel is accretion-induced collapse (AIC) of a massive WD (see, e.g., Miyaji1980; Canal1990; Nomoto1991; Freire2014; Wangbo2018). In this scenario, if the compact object in an IMXB is a massive oxygen-neon-magnesium (ONeMg) WD (close to the Chandrasekhar mass, $M_{\rm Ch}$), accretion onto the WD may supply sufficient material for its mass to exceed $M_{\rm Ch}$. However, rapid rotation might prevent a super-Chandrasekhar WD from collapsing into a NS (Tauris2013aic; Freire2014). After the subsequent CE phase, the WD may lose angular momentum due to r-mode instability and/or magnetic dipole radiation, potentially leading to collapse into a MSP (Freire2014). It should be noted that the formation of NS+He star systems through this channel requires rather stringent conditions (a narrow initial white dwarf mass range, specific accretion rate ranges, etc.), thus also introducing significant uncertainty. Further investigation is needed in the future work.
@@
The biggest uncertainty in this work pertains to CE evolution (see Ivanova2013, for a review), which determines the properties of the resulting NS + He binaries. The criteria for CE evolution, which are closely intertwined with MT stability criteria, remain a subject of ongoing debate (Soberman1997; Pavlovskii2015; Ge2020; Shao2021; Deng2024a). The post-CE evolutionary fate of a binary system depends on the relative magnitudes of the envelope binding energy and the orbital energy (Webbink1984; Nelemans2005; Klencki2021). However, significant ambiguity surrounds the choice of the core-envelope boundary for calculating the CE binding energy, and different boundary definitions can significantly impact the calculated value (Dewi2000; Dominik2012; Marchant2021; Klencki2021; Gallegos2023). The CE ejection efficiency, $\alpha_{\rm CE}$, also carries substantial uncertainty. Theoretically, $\alpha_{\rm CE}$ should be less than or equal to unity. Studies of WD binaries suggest the $\alpha_{\rm CE}$ values of 0.20.30.2-0.3 (Zorotovic2010; Davis2012; Nandez2015; Ge2022; Scherbak2022; Zorotovic2022). However, research on the formation of black hole LMXBs suggests that $\alpha_{\rm CE}$ is expected to be significantly greater than unity for successful envelope ejection and binary survival  [^3] Some studies suggest that if BHs form via failed supernovae (O'Connor2011), the formation of BH LMXBs can be explained without invoking unrealistically high $\alpha_{\rm CE}$ values (Wang2016MN; Shao2020; Deng2024b). (Kalogera1999; Podsiadlowski2003; Kiel2008; Yungelson2008). Studies of Galactic DNSs also find that $\alpha_{\rm CE}$ needs to exceed unity to match observations (Chu2022; Sgalletta2023). Some researchers propose that additional, as-yet-unknown energy sources, supplementing the orbital energy, may contribute to envelope removal (Ivanova2002; Podsiadlowski2010). Currently, fully three-dimensional hydrodynamic simulations of the CE phase in NS HMXBs remain computationally prohibitive (e.g., Moreno2022; Vetter2024).
@@ 5 Summary
In this work, we investigate NS accretion in intermediate-mass or high-mass X-ray binaries during the RLO phase preceding CE evolution. Motivated by the discovery of nearly ten NS ULXs (Kuranov2021), we consider super-Eddington accretion-induced field decay and spin-up in the NSs. By combining MESA with the binary population synthesis code BSE, we explore the formation of NS + He star systems and determine the distributions of the orbital period, He star mass, NS spin period, magnetic field strength, and accreted mass. Our main findings are as follows: